OpenStudy (anonymous):

Find the geometric mean of the pair of numbers .. A) 4√2 B) 32 C) 2√10

OpenStudy (anonymous):

Pair? You have 3 numbers.

OpenStudy (anonymous):

Yeah I know but I already know the other one is wrong... The other one is 6

OpenStudy (anonymous):

I don't know what pair but just sqrt of product

OpenStudy (anonymous):

The geometric mean of 2 numbers a and b is sqrt(ab)

OpenStudy (anonymous):

4*8= 32?

hartnn (hartnn):

so two numbers are 4 and 8 ??

OpenStudy (anonymous):

Where are you getting 4 and 8? There is no reference to either in the question. In fact, I'm not sure you are stating the full question.

OpenStudy (anonymous):

Oh , I didnt realize I didnt put them.. sorry

OpenStudy (anonymous):

Its a part of the question

OpenStudy (anonymous):

The geometric mean of 4 and 8 is A.

hartnn (hartnn):

geometric mean is the square root of the product, hence \(\sqrt32=\sqrt{2*4^2}=4\sqrt2\)

OpenStudy (anonymous):

Okay so how would I find the geoemtricmean of the pair of numbers 3 and 4?

OpenStudy (anonymous):

sqrt(3 x 4)

hartnn (hartnn):

square root of product again. \(\sqrt{3*4}=\sqrt{3*2^2}=?\)

OpenStudy (anonymous):

3*4=12

OpenStudy (anonymous):

yes, and then take the square root of that.

OpenStudy (anonymous):

how do you find the square root.. im confused on that part

OpenStudy (anonymous):

Just put the sqrt sign around the product. As for simplifying, sqrt(12) = sqrt(2 x 2 x 3) = 2 x sqrt(3)

OpenStudy (anonymous):

That's all there is to it for this type of question on the geometric mean.

OpenStudy (anonymous):

Im not understanding.. the sqrt is 12.. so 3*12=36 ... 4*12=48 .. would that mean its 2√3

OpenStudy (anonymous):

It looks like you are looking for r. You are looking for the number by which you put an exponent to get to the next term, right?

OpenStudy (anonymous):

I honestly dont know what I'm doing because all the problem says is Find the geometric mean of the pair of numbers - 3 and 4.. A) 2√3 B) √15 C) 12

OpenStudy (anonymous):

Well, the answer is sqrt(12), so maybe the answer key selection has a misprint. C is most likely the choice sqrt(12).

OpenStudy (anonymous):

Actually, its A, Right in front of me!.

OpenStudy (anonymous):

Because, sqrt(12) is 2 x sqrt(3) as discussed above. So, you DO have your answer!

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

sqrt(12) = sqrt(2 x 2 x 3) = 2 sqrt(3). A.

OpenStudy (anonymous):

There are 2 twos inside the square root sign, so you can pull them OUT.

OpenStudy (anonymous):

Because sqrt(2 x 2 x 3) = sqrt(2 x 2) x sqrt(3) = sqrt(2^2) x sqrt(3) = 2 x sqrt(3). That is ALL of the steps. Do you follow now?

OpenStudy (anonymous):

I do.. thanks.

OpenStudy (anonymous):

Good luck to you on all of your studies.